OpenStudy (anonymous):

Solve: x - 3 = 5x + 13

Parth (parthkohli):

Add 3 to both sides. Bring all the constants on one side and isolate x.

OpenStudy (anonymous):

x - 3 = 5x + 13 x = 5x + 13 +3 x = 5x + 16 x-5x = 16 -4x = 16 x = -16/4 x = -4